Consider a series RLC circuit with R = 15 , L = 206 mH, C = 75 µF, and a maximum voltage of 147 V.
(a) What is the impedance of the circuit at resonance?

(b) What is the resonance frequency of the circuit?
1
Expert's answer
2018-09-26T10:22:08-0400
a) The impedance of the circuit at resonance (X_L-X_C=0):
Z=R
Z=15 Ω.
b) The resonance frequency of the circuit:
f=1/(2π√LC)
f=1/(2π√(0.206∙75∙〖10〗^(-6) )) =40.5 Hz